OpenStudy (anonymous):

What are three laws of exponents (for multiplication, division, and powers)? Write and explain each law. Why do we use exponents in scientific notation (you can provide an example)

TheSmartOne (thesmartone):

\(\Huge\color{red}{a^m\times a^n= a^{m+n}}\) \(\Huge\color{blue}{\frac{a^m}{a^n}=a^{m-n}}\) \(\Huge\color{green}{(a^m)^n=a^{m \times n}}\)